OpenStudy (anonymous):

y = 1.5(2.4)x represent growth or decay?

jimthompson5910 (jim_thompson5910):

The basic form is y = a*b^x Looking at y = 1.5*(2.4)^x, what is the value of 'b'?

OpenStudy (anonymous):

2.4?

jimthompson5910 (jim_thompson5910):

is that smaller than 1 or larger than 1?

OpenStudy (anonymous):

There is no exponent there...

OpenStudy (anonymous):

um larger?

jimthompson5910 (jim_thompson5910):

so \[\Large y = 1.5(2.4)^x\] represents exponential growth

jimthompson5910 (jim_thompson5910):

if the value of b was smaller than 1, then you'd have exponential decay

OpenStudy (anonymous):

than its growth

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

Looks like it's just the slope of the line Jim. XHUHX: The basic form of the equation of a line is y=mx+b, where m=slope and b=y-intercept. As long as your m value is larger than 0, it is growing. (positive slope) If m is negative the line is sloping down and therefore decaying. If you are talking about exponential growth and decay, Jim is right. The way to write something to the power of something else (using your question) would be y=1.5(2.4^x).